# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

### v1

@@ 7 Conclusions and outlook
Let $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$ with $\mathcal{B}$ the Borel sigma-algebra on $\mathbb{R_{+}}$. We will solve $\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p)$, which can be rewritten as
@@
Figure 7: All four scenarios with $\mu=1$, $\underset{\bar{}}{\sigma}=0.1$, $\bar{\sigma}=1$, and $\beta=4$.
@@
The point $\alpha$ can be determined by considering a two-point distribution with support $\{p,\alpha\}$, mean $\mu$, and variance $\bar{\sigma}^{2}$, yielding
$\alpha=\mu+\frac{\bar{\sigma}^{2}}{\mu-p}.$
Consequently,
@@
Finally, in scenario 4, we note that $F(x)=0$, for all $x$ in the interval $[0,\beta]$. This gives $\lambda_{0}=0$, $\lambda_{1}=0$, $\lambda_{2}=0$, and $\lambda_{3}=0$. Hence,
@@
holding pointwise for each $x\in[0,\beta]$. Equation (63) in conjunction with equation (64) is known as the complementary slackness relationship in semi-infinite linear programming. A direct implication of this principle is that the atoms of the optimal primal solution coincide with the intersection of $\lambda^{*}_{0}+\lambda^{*}_{1}x+(\lambda^{*}_{2}+\lambda^{*}_{3})x^{2}$ and the indicator function $\mathds{1}\{x>p\}$. Following this, we will present four viable dual solutions corresponding to these scenarios.
@@
$p_{p}=\frac{\beta-\mu}{\beta-p},\quad p_{L}=\frac{\mu-p}{\beta-p},$ (67)
@@
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{L}=\frac{\mu-p}{\beta-p}.$ (68)
However, the feasibility criterion necessitates
$\underaccent{\bar}{\sigma}^{2}\leq(p-\mu)^{2}p_{p}+(\beta-\mu)^{2}p_{L}\leq\bar{\sigma}^{2},$ (69)
requiring $p\in[\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1}]$.
@@
$p_{0}=\frac{(\beta-\mu)(p-\mu)+\underaccent{\bar}{\sigma}^{2}}{Lp},\quad p_{p}=1-p_{0}-p_{L},\quad p_{L}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (70)
@@
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{L}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (71)
However, this distribution is only feasible when $p_{0}\geq 0$ and $p_{L}\geq 0$, which requires $p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$.
@@
However, this distribution is only feasible when $\beta\geq 0$, requiring $p\in[\underaccent{\bar}{\upsilon}_{2},\beta]$. Since the primal and dual objective values are the same, we know that strong duality holds. Therefore, these objective values are optimal. In addition, the objective value is left-continuous in $p$. Therefore, we can apply Lemma A.1, stated below, and conclude that
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p).$ (74)
This completes the proof.
@@
Let $p\in(0,\beta]$, $\mathcal{P}$ any ambiguity set, and assume that
$\lim_{\epsilon\xrightarrow{}p^{-}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>\epsilon)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (75)
Then
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (76)
Proof. Let $p\in(0,\beta]$ and $\epsilon>0$. The following identity trivially holds:
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon).$ (77)
@@
However, since $\epsilon$ is chosen arbitrarily, we also have that
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\lim_{\epsilon\xrightarrow{}0^{+}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon)=\lim_{\epsilon\xrightarrow{}p^{-}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>\epsilon)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ (78)
Hence, we obtain
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p),$ (79)
which proves the claim. ∎
Appendix B Remaining proofs of pricing results
B.1 Proof of Lemma 4.2

### v2

@@ Appendix A Tail bound results
Let $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$ with $\mathcal{B}$ the Borel sigma-algebra on $\mathbb{R_{+}}$. We solve $\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p)$, which can be rewritten as
@@
Figure 16: Visualization of scenario 1 (top-left), scenario 2 (top-right), scenario 3 (bottom-left), and scenario 4 (bottom-right) with $\mu=1$, $\underset{\bar{}}{\sigma}=0.1$, $\bar{\sigma}=1$, and $\beta=4$.
@@
The point $\alpha$ can be determined by considering a two-point distribution with support $\{p,\alpha\}$, mean $\mu$, and variance $\bar{\sigma}^{2}$, yielding $\alpha=\mu+\bar{\sigma}^{2}/(\mu-p)$. Consequently,
@@
Finally, in scenario 4, we note that $F(x)=0$, for all $x$ in the interval $0,\beta]$. This gives $\lambda_{0}=0$, $\lambda_{1}=0$, $\lambda_{2}=0$, and $\lambda_{3}=0$. Hence,
@@
holding point-wise for each $x\in[0,\beta]$. Equation (47) in conjunction with equation (48) is known as the complementary slackness relationship in semi-infinite linear programming. A direct implication of this principle is that the atoms of the optimal primal solution coincide with the intersection of $\lambda^{*}_{0}+\lambda^{*}_{1}x+(\lambda^{*}_{2}+\lambda^{*}_{3})x^{2}$ and the indicator function $\mathds{1}\{x>p\}$. Following this, we will present four viable dual solutions corresponding to these scenarios.
@@
$p_{p}=\frac{\beta-\mu}{\beta-p},\quad p_{\beta}=\frac{\mu-p}{\beta-p},$ (51)
@@
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{\beta}=\frac{\mu-p}{\beta-p}.$ (52)
However, the feasibility criterion necessitates $\underaccent{\bar}{\sigma}^{2}\leq(p-\mu)^{2}p_{p}+(\beta-\mu)^{2}p_{\beta}\leq\bar{\sigma}^{2}$, requiring $p\in[\bar{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{1}]$.
@@
$p_{0}=\frac{(\beta-\mu)(p-\mu)+\underaccent{\bar}{\sigma}^{2}}{\beta p},\quad p_{p}=1-p_{0}-p_{\beta},\quad p_{\beta}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (53)
@@
$\int_{x}\mathds{1}\{x>p\}{\rm d}\mathbb{P}(x)=p_{\beta}=\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}-\mu p}{\beta(\beta-p)}.$ (54)
However, this distribution is only feasible when $p_{0}\geq 0$ and $p_{\beta}\geq 0$, which requires $p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$.
@@
However, this distribution is only feasible when $\beta\geq 0$, requiring $p\in[\underaccent{\bar}{\upsilon}_{2},\beta]$. Since the primal and dual objective values are the same, we know that strong duality holds. Therefore, these objective values are optimal. In addition, the objective value is left-continuous in $p$. Therefore, we can apply Lemma A.1, stated below, and conclude that $\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p).$ This completes the proof.
@@
Let $p\in(0,\beta]$ and $\mathcal{P}$ any ambiguity set.
  1. (i)
    If $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)$ is left-continuous in $p$, then $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$
  2. (ii)
    If $\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)$ is right-continuous in $p$, then $\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)=\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$
Proof. Let $p\in(0,\beta]$ and $\epsilon>0$. The following identity trivially holds: $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon).$ However, since $\epsilon$ is chosen arbitrarily, we also have that $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\lim_{\epsilon\xrightarrow{}0^{+}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p-\epsilon)=\lim_{\epsilon\xrightarrow{}p^{-}}\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>\epsilon)=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p).$ Hence, we obtain $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p),$ which proves (i).
@@
Next, observe that $\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\geq\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\geq\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p+\epsilon).$ However, since $\epsilon$ is chosen arbitrarily, we also have that $\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\geq\lim_{\epsilon\xrightarrow{}0^{+}}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p+\epsilon)=\lim_{\epsilon\xrightarrow{}p^{+}}\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq\epsilon)=\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p).$ Hence, we obtain $\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p)\leq\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p)\leq\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X\geq p),$ which proves (ii). ∎
A.2 Tight upper tail bounds
Lemma A.2.
Consider ambiguity set $\mathcal{P}_{(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}$ and denote $\underaccent{\bar}{\upsilon}_{1}:=\mu-\frac{\underaccent{\bar}{\sigma}^{2}}{\beta-\mu}$, $\underaccent{\bar}{\upsilon}_{2}:=\mu+\frac{\underaccent{\bar}{\sigma}^{2}}{\mu}$ and $\bar{\upsilon}_{2}:=\mu+\frac{\bar{\sigma}^{2}}{\mu}$. Then,
$\sup_{\mathbb{P}\in\mathcal{P}_{(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}}\mathbb{P}(X\geq p)=\sup_{\mathbb{P}\in\mathcal{P}_{(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}}\mathbb{P}(X>p)=\begin{cases}1,\quad&p\in(0,\underaccent{\bar}{\upsilon}_{1}],\\ \frac{(\beta+p)\mu-\mu^{2}-\underaccent{\bar}{\sigma}^{2}}{\beta p},\quad&p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}],\\ \frac{\mu}{p},\quad&p\in[\underaccent{\bar}{\upsilon}_{2},\bar{\upsilon}_{2}],\\ \frac{\bar{\sigma}^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}},&p\in[\bar{\upsilon}_{2},\beta].\end{cases}$ (57)
Proof. Let $\mathcal{M}$ be the set of probability measures defined on the measurable space $([0,\beta],\mathcal{B})$. We will solve $\sup_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)$, which can be rewritten as
$\!\sup_{\mathbb{P}\in\mathcal{M}}$ $\int_{x}\mathds{1}_{\{x\geq p\}}{\rm d}\mathbb{P}(x),$ (58)
Consider the corresponding dual,
$\inf_{\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3}}$ $\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2}),$ (59)
s.t. $\mathds{1}_{\{x\geq p\}}\leq\lambda_{0}+\lambda_{1}x+(\lambda_{2}+\lambda_{3})x^{2}=:F(x),\ \forall x\in[0,\beta],\lambda_{2}\leq 0,\lambda_{3}\geq 0.$
The proof is structured analogously to the proof of Lemma 3.1. Hence, we derive both feasible dual and primal solutions and show that strong duality holds through matching objective values. Four scenarios are considered. In the first scenario, $F(x)$ intersects with $\mathds{1}\{x\geq p\}$ within the interval $p,\beta]$. In the second scenario, the intersection occurs in the points $\{0,p,\beta\}$. For the third scenario, the points of intersection are identified at $\{0,p\}$. Finally, in the fourth scenario, the intersections are at $\{\alpha,p\}$, where $\alpha\in[0,p]$. We now present four viable dual solutions corresponding to these scenarios.
For scenario 1 we observe that $F(x)=1$ for all $x$ in the interval $0,\beta]$. This gives $\lambda_{0}=1,\lambda_{1}=0,\lambda_{2}=0$, and $\lambda_{3}=0$. Hence,
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=1.$ (60)
In scenario 2, the conditions $F(0)=0$, $F(p)=1$, and $F(\beta)=1$ are observed. Furthermore, this scenario is characterized by $F$ taking the form of a downward-sloping parabola. To align with these conditions, we set $\lambda_{3}=0$, as $\lambda_{2}\leq 0$. This gives
$\lambda_{0}=0,\ \lambda_{1}=\frac{1}{\beta}+\frac{1}{p},\ \lambda_{2}=-\frac{1}{\beta p},\ \lambda_{3}=0,\$(61)
and
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{(\beta+p)\mu-\mu^{2}-\underaccent{\bar}{\sigma}^{2}}{\beta p}.$ (62)
In scenario 3 we observe that $F(0)=0$ and $F(p)=1$. Moreover, $F$ is linear. Hence, we set $\lambda_{0}=\lambda_{2}=\lambda_{3}=0$, resulting in
$\lambda_{0}=0,\ \lambda_{1}=\frac{1}{p},\ \lambda_{2}=0,\ \lambda_{3}=0,\$(63)
and
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{\mu}{p}.$ (64)
Finally, in scenario 4, we note that $F(p)=1$, $F(\alpha)=0$, and $F^{\prime}(\alpha)=0$. In this scenario, the functional form of $F$ is an upward-sloping parabola. Hence, following these specified conditions, we set $\lambda_{2}=0$, as $\lambda_{3}\geq 0$. This gives dual solution
$\lambda_{0}=\frac{\alpha^{2}}{(\alpha-p)^{2}},\ \lambda_{1}=-\frac{2\alpha}{(\alpha-p)^{2}},\ \lambda_{2}=0,\ \lambda_{3}=\frac{1}{(\alpha-p)^{2}}.$ (65)
The point $\alpha$ can be determined by considering a two-point distribution with support $\{\alpha,p\}$, mean $\mu$, and maximal variance $\bar{\sigma}^{2}$, giving
$\alpha=\mu-\frac{\bar{\sigma}^{2}}{p-\mu},$
and
$\lambda_{0}+\lambda_{1}\mu+\lambda_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda_{3}(\bar{\sigma}^{2}+\mu^{2})=\frac{\bar{\sigma}^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}}.$ (66)
$\int_{x}\mathds{1}\{x\geq p\}{\rm d}\mathbb{P}^{*}(x)=\int_{x}(\lambda^{*}_{0}+\lambda^{*}_{1}\mu+\lambda^{*}_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda^{*}_{3}(\bar{\sigma}^{2}+\mu^{2})){\rm d}\mathbb{P}^{*}(x).$ (67)
Furthermore, due to dual feasibility, we obtain
$\lambda^{*}_{0}+\lambda^{*}_{1}\mu+\lambda^{*}_{2}(\underaccent{\bar}{\sigma}^{2}+\mu^{2})+\lambda^{*}_{3}(\bar{\sigma}^{2}+\mu^{2})-\mathds{1}\{x\geq p\}\geq 0,$ (68)
holding point-wise for each $x\in[0,\beta]$. Equation (67) in conjunction with equation (68) is known as the complementary slackness relationship in semi-infinite linear programming. A direct implication of this principle is that the atoms of the optimal primal solution coincide with the intersection of $\lambda^{*}_{0}+\lambda^{*}_{1}x+(\lambda^{*}_{2}+\lambda^{*}_{3})x^{2}$ and the indicator function $\mathds{1}\{x\geq p\}$.
In scenario 1, the construction involves a two-point distribution with atoms $\{p,\bar{\alpha}\}$, where $\bar{\alpha}:=\mu+\frac{\underaccent{\bar}{\sigma}^{2}}{\mu-p}$. The corresponding probabilities are
$p_{p}=\frac{\underaccent{\bar}{\sigma}^{2}}{(p-\mu)^{2}+\underaccent{\bar}{\sigma}^{2}},\quad p_{\bar{\alpha}}=1-p_{p},$ (69)
leading to
$\int_{x}\mathds{1}\{x\geq p\}{\rm d}\mathbb{P}(x)=1.$ (70)
However, this distribution’s feasibility requires $\bar{\alpha}\leq\beta$, necessitating $p\in[0,\underaccent{\bar}{\upsilon}_{1}]$.
For scenario 2, a unique three-point distribution is considered with probability masses at $\{0,p,\beta\}$. The probabilities are
$p_{0}=1-\frac{(\beta+p)\mu-\mu^{2}-\underaccent{\bar}{\sigma}^{2}}{\beta p},\quad p_{p}=1-p_{0}-p_{\beta},\quad p_{\beta}=\frac{\underaccent{\bar}{\sigma}^{2}+\mu^{2}-p\mu}{\beta(\beta-p)},$ (71)
resulting in
$\int_{x}\mathds{1}\{x\geq p\}{\rm d}\mathbb{P}(x)=p_{p}+p_{\beta}=\frac{(\beta+p)\mu-\mu^{2}-\underaccent{\bar}{\sigma}^{2}}{Lp}.$ (72)
However, this distribution is only feasible when $p_{0}\geq 0$ and $p_{\beta}\geq 0$, requiring $p\in[\underaccent{\bar}{\upsilon}_{1},\underaccent{\bar}{\upsilon}_{2}]$.
For scenario 3, we consider a two-point distribution on $\{0,p\}$ with probabilities
$p_{0}=1-p_{p},\quad p_{p}=\frac{\mu}{p}.$ (73)
This yields
$\int_{x}\mathds{1}\{x\geq p\}{\rm d}\mathbb{P}(x)=p_{p}=\frac{\mu}{p}.$ (74)
The feasibility criterion necessitates
$\underaccent{\bar}{\sigma}^{2}\leq\mu^{2}p_{0}+(p-\mu)^{2}p_{p}\leq\bar{\sigma}^{2},$ (75)
requiring $p\in[\underaccent{\bar}{\upsilon}_{2},\bar{\upsilon}_{2}]$.
Scenario 4 sees a two-point distribution with probability masses on the points $\{\alpha,p\}$ with probabilities
$p_{\alpha}=1-p_{p},\quad p_{p}=\frac{\bar{\sigma}^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}},$ (76)
leading to
$\int_{x}\mathds{1}\{x\geq p\}{\rm d}\mathbb{P}(x)=p_{p}=\frac{\bar{\sigma}^{2}}{(\mu-p)^{2}+\bar{\sigma}^{2}}.$ (77)
However, this distribution is only feasible when $\alpha\leq\beta$, requiring $p\in[\bar{\upsilon}_{2},\beta]$. Since the primal and dual objective values are the same, we know that strong duality holds. Therefore, these objective values are optimal. Furthermore, the objective value is right-continuous in $p$. Therefore, we can apply Lemma A.1 and conclude that $\sup_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X\geq p)=\sup_{\mathbb{P}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)}\mathbb{P}(X>p).$
Appendix B Comparison with optimal pricing
As is common in the pricing literature, we will now evaluate the performance of the robust pricing strategy with the optimal Bayesian strategy; see e.g. Azar and Micali (2012), Chen et al. (2022a). The optimal Bayesian strategy is known to be fixed pricing (Riley and Zeckhauser 1983) and achieves the optimal expected revenue $\textup{OPT}(\mathbb{P})\mathrel{\mathop{:}}=\sup_{p}p\mathbb{P}(X\geq p)$. Markov’s inequality yields $p\mathbb{P}(X\geq p)\leq\mu$, and hence $\textup{OPT}(\mathbb{P})\leq\mu$. This is in fact a tight bound. To see this, let $p\in[\mu+{\underaccent{\bar}{\sigma}^{2}}/{\mu},\mu+{\bar{\sigma}^{2}}/{\mu}]$ and consider the two-point distribution $\mathbb{P}_{0}\in\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$ with atoms $\{0,p\}$ and associated probabilities $1-\mu/p$ and ${\mu}/{p}$, respectively, for which $p\mathbb{P}_{0}(X\geq p)=\mu$. Combining this bound with Theorem 3.1 yields, upon some rewriting, the following performance guarantee for robust pricing:
Corollary B.1.
$\frac{p^{*}\mathbb{P}(X\geq p^{*})}{\textup{OPT}(\mathbb{P})}\geq\begin{cases}\rho_{l}(\mu,\bar{\sigma})\mathrel{\mathop{:}}=\frac{1}{2}\frac{\bar{\sigma}}{\mu}\kappa^{3}(\mu,\bar{\sigma}),\quad&(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{l},\\ \rho_{m}(\mu,\beta)\mathrel{\mathop{:}}=\frac{2}{\mu}(\beta-\sqrt{\beta(\beta-\mu)})-1,\quad&(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{m},\\ \rho_{h}(\mu,\underaccent{\bar}{\sigma},\beta)\mathrel{\mathop{:}}=2(1-\sqrt{1-\frac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta}})-\tfrac{\mu^{2}+\underaccent{\bar}{\sigma}^{2}}{\mu\beta},\quad&(\underaccent{\bar}{\sigma},\bar{\sigma})\in\Sigma_{h}.\end{cases}$ (78)
[…]
Figure 17: Corollary B.1 with $\mu=0.5$ and $\beta=1$.
U-shaped performance guarantee. Figure 17 illustrates Corollary B.1. When the seller has precise information about the variance, there occurs a U-shaped performance guarantee, aligned with the U-shaped revenue curves found in monopolistic markets (Johnson and Myatt 2006, Sun 2012) discussed in the introduction. In our case, this U-shape occurs as the seller switches from low to high pricing, changing the curve from being decreasing to increasing. More generally, we have found that the U-shape occurs when the seller has precise enough (not perfect per se) information on the variance. When the seller only knows an upper bound, the U-shape is replaced with a function that first decreases and then remains constant, see Figure 17. We substantiate these numerical insights with the following result:
Corollary B.2 (Properties of price and revenue).
Risk-aligned pricing. The optimal price and performance guarantee adjust in the same direction in response to changes in mean, support, and variance. In uncertain markets, robust pricing involves higher prices to reflect risk levels. Higher volatility leads to a high-price strategy, targeting a smaller, high-value customer segment. Corollary B.2 shows that only the high price is positively correlated with variance, resulting in a performance guarantee that increases with variance. Managers in volatile environments should focus on value alignment rather than broad appeal, catering to customers willing to pay for exclusivity, as seen in luxury fashion and high-end tech. Conversely, Corollary B.2 shows that low prices are negatively correlated with variance. For example, electric vehicles initially targeted a niche market with high prices due to production and demand uncertainties. As technology matured and demand increased, prices dropped, attracting mainstream consumers and a larger customer base.
Appendix C Remaining proofs of pricing results
C.1 Proof of Lemma 4.2